# Benchmarking Deep Learning-Based Methods for Irradiance Nowcasting with Sky Images

### v1

@@ Abstract
To address the high levels of uncertainty associated with photovoltaic energy, an increasing number of studies focusing on short-term solar forecasting have been published. Most of these studies use deep learning-based models to directly forecast a solar irradiance or photovoltaic power value given an input of sky image sequences. Recently, however, advances in generative modeling have led to approaches that divide the forecasting problem into two sub-problems: 1) future event prediction, i.e. generating future sky images; and 2) solar irradiance or photovoltaic power nowcasting, i.e. predicting the concurrent value from a single image. One such approach is the SkyGPT model, where they show that the potential for improvement is much larger for the nowcasting model than for the generative model. Thus, in this paper, we focus on the solar irradiance nowcasting problem and conduct an extensive benchmark of deep learning architectures across the widely-used Folsom, SIRTA and NREL datasets. Moreover, we perform ablation experiments on different training configurations and data processing techniques, including the choice of the target variable used for training and adjustments of the timestamp alignment between images and irradiance measurements. In particular, we draw attention to a potential error associated with the sky image timestamps in the Folsom dataset and a possible fix is discussed. All our results are reported in terms of both the root mean squared error and the mean absolute error and, by leveraging the three datasets, we demonstrate that our findings are consistent across different solar stations.
@@
Solar Irradiance , Nowcasting , Forecasting , Deep Learning , Sky Images , Computer Vision
@@
[label1, label2]organization=Department of Electrical Engineering, Federal University of Santa Catarina,addressline=, city=Florianópolis, postcode=, state=Santa Catarina, country=Brazil
{highlights}
Solar irradiance nowcasting using all sky images (ASI).
Benchmarking deep learning architectures for the irradiance nowcasting task.
A novel method for sky image and irradiance timestamp alignment is proposed.
Different model target variables are thoroughly investigated.
Findings proved consistent across three widely-used solar irradiance datasets.
@@ 1 Introduction
In the very short term solar forecasting ($\leq$ 60 minutes forecasting horizon), all sky image (ASI) based models have achieved state of the art results, due to their ability to capture the cloud information with a very high spatio-temporal resolution [8]. These images are typically combined with convolutional neural networks (CNN), which extract features from the images that will be used to predict the solar value. The following paragraphs briefly describe some of these approaches.
@@
The authors of [9] use a ResNet18 architecture [10] to encode the image features, which are then used to predict the solar irradiance 5 to 10 minutes ahead. However, instead of feeding a static image to the model, they stack multiple images at different timestamps and gain up to a 7% performance boost. The method proposed in [11] used a 3D CNN [12] to forecast the clear sky index $k_{t}$, which results in a 15.2% improvement over the smart persistence model. In [13], a benchmark study of four different architectures (CNN, CNN+LSTM, 3D-CNN and ConvLSTM) is made and, while the more complex models tend to perform better, it is shown that they all share a fundamental problem: the prediction of future irradiance values often lag behind the ground truth measurements. A multi-modal transformer based framework is proposed in [14], where the solar irradiance time series is encoded by a transformer [15], and the sky images are transformed into optical flow maps and encoded by a vision transformer [16], which are then fused with a cross modality attention block to predict the solar irradiance up to 30 minutes ahead. In [17], a VGG16 [18] inspired model is developed specifically for the solar irradiance forecasting task, the SolarNet model, which outperformed several of their benchmark models. In [19], different input and output configurations are explored for the Sunset model, which is another model specifically designed for the solar forecasting task, but focuses on the solar PV power instead of solar irradiance.
@@
The methodology proposed in [20] is particularly interesting as it separates the future event prediction task from the irradiance mapping task. This is achieved by using a sky image generation model, SkyGPT, which is trained auto-regressively to predict future sky images based on past image sequences. With the generated future sky image, a modified version of the U-Net architecture [21] is then used to predict the concurrent PV power. By decoupling the future event prediction task (SkyGPT) from the PV power nowcasting task (U-Net), they demonstrate that improving the SkyGPT model (such that it could perfectly predict the future sky image) would only boost the performance by 13%, whereas improving the nowcasting model could yield up to a 64% a performance boost (see Figure 7 of [20]).
@@
Inspired by the larger potential improvement, this study focuses primarily on the solar irradiance nowcasting task, using exclusively sky images as input, i.e. given an image at timestamp $t$, we want to predict the corresponding irradiance value at the same timestamp $t$. Although this is different from PV power nowcasting, we expect the results to generalize well across both tasks, since they share the same underlying nature and are highly correlated.
@@
Compared to solar forecasting, fewer studies have tackled the solar nowcasting task. In [22] they perform solar irradiance nowcasting and show that appending a Sun mask as a fourth channel to the images significantly increases the nowcasting performance, which we also investigate in this study. In [23] they adapt the Sunset model for the solar PV power nowcasting task, which is used in this study for the irradiance nowcasting.
@@
We begin our study by conducting an extensive benchmark of ten different deep learning architectures across the Folsom, SIRTA and NREL datasets, which are three datasets commonly used in the solar forecasting literature. We then perform several ablation experiments, including the choice of the target variable used for training and adjusting the timestamp alignment between the images and irradiance measurements. All of our experiments are evaluated using the root mean squared error (RMSE) and mean absolute error (MAE) and we show that our findings are consistent across all three datasets. Our main contributions can be summarized as follows:
@@
    We conduct an extensive benchmark of deep learning architectures across three solar irradiance datasets and find that the ResNet50 architecture provides significant gains over the U-Net architecture used in [20];
@@ 2.1 Irradiance Components
[…]
@@
where $\theta_{z}$ is the zenith angle. This component is the primary focus of this study.
@@ 2.2 Clear Sky Models
Predicting the GHI under clear sky conditions is a well-defined problem, and there are hundreds of different clear sky models available in the literature, with varying but generally excellent performances across many different climates [25, 26]. Clear sky models can also be used to obtain the clear sky index $k_{t}$, that normalizes the GHI measurement according to Equation 2:
@@
Other than normalization, the $k_{t}$ index and the clear sky prediction $I^{\text{clr}}$ can be used to obtain the smart persistence model, which is a naive model that assumes that the current $k_{t}$ will stay the same throughout the entire forecasting horizon $h$, as shown in Equation 3. This model is commonly used as a baseline reference for solar forecasting models.
@@
Another commonly used index is the clearness index $K_{t}$ [29], which uses the extraterrestrial irradiance instead to normalize the GHI, as depicted in Equation 5
@@
where $I_{0}$ is the solar constant and $\theta_{z}$ is the zenith angle. The choice between both normalization indexes is investigated in subsubsection 5.2.3.
@@ 3 Data Description and Processing
In this section we describe the raw data used in this study, obtained from three different datasets, as well as the preprocessing pipeline that is made prior to training the models.
@@ 3.1 Folsom Dataset
Examples of the sky images are shown in the leftmost column of Figure 2 and, unlike the other two datasets, in $\mathcal{D}^{\text{folsom}}$ the camera exposure time is the same for all of the images. The GHI distribution of $\mathcal{D}^{\text{folsom}}$ is shown in the left panel of Figure 3, where it is clear that the density of large irradiance values in this dataset is much higher, i.e. it has a larger number of clear sky instants than the other two datasets.
@@ 3.2 SIRTA Dataset
The sky images were captured with a $768\times 1024$ resolution at a 1-minute sampling interval in 2017 and 2-minute in 2018 and 2019 and come in two forms: 1) a long exposure form that shows the sky in more detail but can have saturated pixels in the area around the Sun; 2) a short exposure form that shows the area around the Sun less saturated and in more detail, but the rest of the sky is shown in less detail (middle images of Figure 2). The GHI and DHI are measured by a CMP22 pyranometer and the DNI is measured by a CHP1 pyrheliometer, all at a 1-second sampling interval, but only the 1-minute averages are provided.
@@ 3.3 NREL Dataset
Lastly, we also analyzed the National Renewable Energy Laboratory (NREL) dataset ($\mathcal{D}^{\text{nrel}}$) [46], collected at NREL’s Solar Radiation Research Laboratory (SRRL), located in Golden, Colorado (39.742 N and 105.18 W). The SSRL has been measuring irradiance correlated data since 1981 and, once again, many studies have been published using this dataset: [47, 48, 14, 49, 50]. For our benchmark studies, we only selected the sky images and irradiance data collected from 2020 to 2023. It is important to note that $\mathcal{D}^{\text{nrel}}$ provides multiple measurements for the same type of variable, for the sake of redundancy. In our case, the irradiance data used was:
@@
    GHI: computed from the two above and the solar zenith angle $\theta_{z}$ as per Equation 1.
Like in the previous two datasets, the irradiance data in $\mathcal{D}^{\text{nrel}}$ are 1-minute averages and their GHI distribution is plotted in the right panel of Figure 3.
The sky images are provided at a $1536\times 1536$ pixel resolution in both long and short exposures (rightmost images of Figure 2), however they are only available once every 10 minutes, which makes this a much smaller dataset compared to the other two (for our selected years in each dataset). The impact that a smaller training dataset has on the model performance is further discussed in subsubsection 5.2.5
Figure 2: Sky image examples for each dataset used in this study. $\mathcal{D}^{\text{sirta}}$ and $\mathcal{D}^{\text{nrel}}$ also provide a short exposure alternative for each image, so that the area around the Sun can be shown in greater detail. The depicted images were cropped such that they have the same shape.
Figure 3: The GHI data distributions for the three datasets used in this study, plotted after the data processing pipeline.
@@ 3.4 Sky Image Data Processing
The image processing was essentially the same for all 3 datasets. First we compute the mean image of each dataset (using only the long exposure images for $\mathcal{D}^{\text{sirta}}$ and $\mathcal{D}^{\text{nrel}}$) and then define a circular region of interest (ROI), with radius $R$ and center $C$, which is used as a binary mask for that dataset. These masks are applied to all the images in their respective datasets to cancel every pixel outside their ROI. Following this, we center crop the images on $C$ with length $2R$. Finally, the images are downscaled to a $64\times 64$ resolution using the Python’s PIL library. To avoid any potential JPEG compression losses [51], the resized images are saved to disk in .npy format.
@@ 3.5 Irradiance Data Processing
The irradiance data processing was also largely the same for all 3 datasets. First, we resample the data to a 1-second frequency through linear interpolation (more details on why this is done in the following section). To avoid large interpolation errors, this is only done for the samples that are 1-minute apart, i.e. periods with 2 or more minutes of missing data are discarded. Following the linear interpolation, we remove all samples where $\theta_{z}>80^{\circ}$, as these samples provide little information and the GHI is always low.
The next processing step is only applied to the training set of each dataset (see subsection 3.8 for the train/test splits) and consists of shifting the irradiance measurements in time by applying a time offset ($\Delta\text{t}$), which we do according to (6)
@@
where $T_{\text{new}}$ and $T_{\text{orig.}}$ are the new and original irradiance measurement timestamps, respectively. If a $\Delta\text{t}<0$ is applied, then we are advancing the irradiance series in time, relative to the sky image series, whereas a $\Delta\text{t}>0$ would have the opposite effect, as shown in Figure 4. Depending on how these irradiance values were averaged, we hypothesize that this can be beneficial. For instance, if the value of $I$ at 14:00:00 represents the average GHI values from 13:59:01 to 14:00:00 (backward average), then a $\Delta\text{t}<0$ could diminish the delay introduced by such an average. However, applying a $\Delta\text{t}$ to the testing set would result in different labels for every searched $\Delta\text{t}$, possibly leading to an unfair evaluation. Thus, we only shift the irradiance measurements in the training set, by -20s, 10s and -20s in $\mathcal{D}^{\text{folsom}}$, $\mathcal{D}^{\text{sirta}}$ and $\mathcal{D}^{\text{nrel}}$, respectively. Other time offsets are investigated in subsubsection 5.2.2.

### v2

# Benchmarking Deep Learning Methods for Irradiance Estimation from Sky Images with Applications to Video Prediction-Based Irradiance Nowcasting

@@ Abstract
To address the high levels of uncertainty associated with photovoltaic energy, an increasing number of studies focusing on short-term solar forecasting (i.e. nowcasting) have been published. Most of these studies use deep-learning-based models to directly forecast a solar irradiance or photovoltaic power value given an input sequence of sky images. Recently, however, advances in generative modeling have led to approaches that divide the nowcasting problem into two sub-problems: 1) future event prediction, i.e. generating future sky images; and 2) solar irradiance or photovoltaic power estimation, i.e. predicting the concurrent value from a single image. One such approach is the SkyGPT model, whose potential for improvement is shown to be much larger in the estimation component than in the generative component. Thus, in this paper, we focus on the solar irradiance estimation problem and conduct an extensive benchmark of deep learning architectures across the widely-used Folsom, SIRTA and NREL datasets. Moreover, we perform ablation experiments on different training configurations and data processing techniques, including the choice of the target variable used for training and adjustments of the timestamp alignment between images and irradiance measurements. In particular, we draw attention to a potential error associated with the sky image timestamps in the Folsom dataset and suggest a possible fix. By leveraging the three datasets, we demonstrate that our findings are consistent across different solar stations. Finally, we combine our best irradiance estimation model with a video prediction model and obtain state-of-the-art results on the SIRTA dataset.
@@
Solar Irradiance , Nowcasting , Forecasting , Deep Learning , Sky Images , Video Prediction
@@
[label1, label2]organization=Department of Electrical and Electronic Engineering, Federal University of Santa Catarina,addressline=, city=Florianópolis, postcode=, state=Santa Catarina, country=Brazil
@@ 1 Introduction
In the very short term solar forecasting ($\leq$ 60 minutes forecasting horizon), commonly referred to as solar nowcasting, all-sky image (ASI) based models have achieved state-of-the-art results, due to their ability to capture the cloud information with a very high spatio-temporal resolution [8]. These images are typically combined with convolutional neural networks (CNN), which extract features from the images that will be used to predict the solar value. The following paragraphs briefly describe some of these approaches.
@@
The authors of [9] use a ResNet18 architecture [10] to encode the image features, which are then used to predict the solar irradiance 5 to 10 minutes ahead. However, instead of feeding a static image to the model, they stack multiple images at different timestamps and gain up to a 7% performance boost. The method proposed in [11] used a 3D CNN [12] to forecast the clear sky index $k_{t}$, which results in a 15.2% improvement over the smart persistence model. In [13], a benchmark study of four different architectures (CNN, CNN+LSTM, 3D-CNN and ConvLSTM) is conducted and, while the more complex models tend to perform better, it is shown that they all share a fundamental problem: the predicted future irradiance values are often delayed relative to the ground truth measurements, resulting in irradiance ramp predictions that are systematically delayed relative to the ground truth.
A multi-modal transformer based framework is proposed in [14], where the solar irradiance time series is encoded by a transformer [15], and the sky images are transformed into optical flow maps and encoded by a vision transformer [16], which are then fused with a cross modality attention block to predict the solar irradiance up to 30 minutes ahead. In [17], a VGG16 [18] inspired model is developed specifically for the solar irradiance nowcasting task, the SolarNet model, which outperformed several of their benchmark models. In [19], different input and output configurations are explored for the Sunset model, which is another model specifically designed for the solar nowcasting task, but focuses on the solar PV power instead of solar irradiance.
@@
The methodology proposed in [20] is particularly interesting as it separates the future event prediction task from the irradiance estimation task. This is achieved by using a sky image generation model, SkyGPT, which is trained auto-regressively to predict future sky images based on past image sequences. With the generated future sky image, a modified version of the U-Net architecture [21] is then used to estimate the concurrent PV power. By decoupling the future event prediction task (SkyGPT) from the PV power estimation task (U-Net), they demonstrate that improving the SkyGPT model (such that it could perfectly predict the future sky image) would only boost the performance by 13%, whereas improving the estimation model could yield up to a 64% performance boost (see Figure 7 of [20]).
@@
Inspired by the larger potential improvement, this study focuses primarily on the solar irradiance estimation task, using exclusively sky images as input, i.e. given an image at timestamp $t$, we want to predict the corresponding irradiance value at the same timestamp $t$. Although this is different from PV power estimation, we expect the results to generalize well across both tasks, since they share the same underlying nature and are highly correlated.
@@
Compared to solar nowcasting, fewer studies have tackled the solar estimation task. In [22] they perform solar irradiance estimation and show that appending a Sun mask as a fourth channel to the images significantly increases the estimation performance, which we also investigate in this study. In [23] they adapt the Sunset model for the solar PV power estimation task, which we use in this study for irradiance estimation.
@@
We begin our study by conducting an extensive irradiance estimation benchmark of ten deep learning architectures across the Folsom, SIRTA and NREL datasets, which are three datasets commonly used in the solar forecasting literature. We then perform several ablation experiments, including the choice of the target variable used for training and adjusting the timestamp alignment between the images and irradiance measurements. Finally, we train a video prediction model and combine it with our irradiance estimation model to conduct proof of concept experiments on the two-step approach in solar irradiance nowcasting. All of our experiments are evaluated using the root mean squared error (RMSE) and mean absolute error (MAE) and we show that our findings are consistent across all three datasets. Our main contributions can be summarized as follows:
@@
    We conduct an extensive irradiance estimation benchmark of deep learning architectures across three solar irradiance datasets and find that the ResNet50 architecture provides significant gains over the U-Net architecture used in [20];
@@
  5. 5.
    With our best irradiance estimation model and a simple video prediction model, we obtain state of art nowcasting results on the SIRTA dataset.
@@ 2.1 Irradiance Components
Solar irradiance is a measurement of the rate of solar radiation energy that arrives at a surface area and is typically measured in $\text{W/m}^{2}$. The mean solar irradiance at the top of the Earth’s atmosphere, also known as the solar constant [24], is approximately 1366 $\text{W/m}^{2}$. After entering the Earth’s atmosphere, the solar radiation will be attenuated by the complex interactions with the atmospheric constituents, resulting in two irradiance components at the ground level: diffuse horizontal irradiance (DHI) and direct normal irradiance (DNI). The DHI is the amount of solar radiation incident upon a horizontal plane that has been scattered or reflected by atmospheric constituents or nearby objects. The DNI measures the amount of solar radiation that is coming directly from the direction of the Sun beam. The sum of these two components on a horizontal plane results in the global horizontal irradiance (GHI), denoted by:
@@
where $\theta_{z}$ is the Solar Zenith Angle (SZA). This component is the primary focus of this study.
@@ 2.2 Clear Sky Models
Predicting the GHI under clear sky conditions is a well-defined problem, and there are hundreds of different clear sky models available in the literature, with varying but generally excellent performances across many different climates [25, 26]. Clear sky models can also be used to obtain the clear sky index $k_{t}$, that normalizes the GHI measurement according to:
@@
Other than normalization, the $k_{t}$ index and the clear sky prediction $I^{\text{clr}}$ can be used to obtain the smart persistence model (SPM), which is a naive model that assumes that the current $k_{t}$ will stay the same throughout the entire forecasting horizon $h$:
@@
This model is commonly used as a baseline reference for solar forecasting models.
Another commonly used index is the clearness index $K_{t}$ [29], which uses the extraterrestrial irradiance instead to normalize the GHI:
@@
where $I_{0}$ is the solar constant and $\theta_{z}$ is the SZA. The choice between both normalization indices is investigated in subsubsection 5.2.3.
@@ 3 Data Description and Processing
In this section, we describe the raw data used in this study, obtained from three different datasets, as well as the preprocessing pipeline that is made prior to training the models.
@@ 3.1 Folsom Dataset
Examples of the sky images are shown in the leftmost column of Figure 2 and, unlike the other two datasets, the camera exposure time is the same for all images in $\mathcal{D}^{\text{folsom}}$, though the exact value of this setting is unavailable in [32] and the image metadata. Also unlike the other datasets, $\mathcal{D}^{\text{folsom}}$ exhibits a few instances of lens contamination due to dust, dirt, and condensation. Although we believe that automatic mitigation strategies for this issue are beyond the scope of this study, we think it is important to highlight it and to inform future research, since this can have an impact on model performance. Thus, we manually identified three representative examples in Figure 3. The GHI distribution of $\mathcal{D}^{\text{folsom}}$ is shown in the left panel of Figure 4, where it is clear that the density of large irradiance values in this dataset is much higher, i.e. it has a larger number of clear sky instances than the other two datasets.
@@ 3.2 SIRTA Dataset
The sky images (middle images of Figure 2) were captured with a $768\times 1024$ resolution at a 1-minute sampling interval in 2017 and 2-minute in 2018 and 2019 and come in two forms: 1) a long exposure form that shows the sky in more detail but can have saturated pixels in the area around the Sun (1/1000 s exposure); 2) a short exposure form that shows the area around the Sun less saturated and in more detail, but the rest of the sky is shown in less detail (1/2000 s exposure). The GHI and DHI are measured by a CMP22 pyranometer and the DNI is measured by a CHP1 pyrheliometer, all at a 1-second sampling interval, but only the 1-minute averages are provided.
@@ 3.3 NREL Dataset
Lastly, we also analyzed the National Renewable Energy Laboratory (NREL) dataset ($\mathcal{D}^{\text{nrel}}$) [46], collected at NREL’s Solar Radiation Research Laboratory (SRRL), located in Golden, Colorado (39.742 N and 105.18 W). The SSRL has been measuring irradiance correlated data since 1981 and, once again, many studies have been published using this dataset: [47, 48, 14, 49, 50]. For our benchmark studies, we only selected the sky images and irradiance data collected from 2020 to 2023. Unlike the other datasets, $\mathcal{D}^{\text{nrel}}$ provides multiple independent measurements of the same irradiance component. We used the following sensors:
@@
    GHI: CMP22, CMP22-1, CMP22-2, CMP11 and CM6b pyranometers.
By leveraging all of these sensors, we were able to detect and filter out several faulty measurements. To do so, we implement the following: let $M(X)$ denote the median of a set of measurements $X$. We first compute the medians of the DHI and DNI measurements, $M(\text{DHI})$ and $M(\text{DNI})$, and use them to estimate the global component, $\text{GHI}_{\text{calculated}}$, according to Equation 1. We then compare this calculated value with the median of the GHI measurements, $M(\text{GHI})$, and keep only the samples that satisfy:
$|\text{GHI}_{\text{calculated}}-M(\text{GHI})|\leq 10\text{ W/m}^{2}$ (6)
This filtering step retains 961,069 of the original 1,063,282 measurements and improves the overall reliability of the irradiance labels. The GHI distribution is plotted in the right panel of Figure 4.
The sky images are provided at a $1536\times 1536$ pixel resolution in both long and short exposure times (rightmost images of Figure 2), though the exact value of this setting is also unavailable in this dataset. The sky images are only available once every 10 minutes, which makes this a much smaller dataset compared to the other two (for our selected years in each dataset). The impact that a smaller training dataset has on the model performance is further discussed in subsubsection 5.2.5
Figure 2: Clear and cloudy sky image examples for each dataset used in this study. $\mathcal{D}^{\text{sirta}}$ and $\mathcal{D}^{\text{nrel}}$ also provide a short exposure alternative for each image, so that the area around the Sun can be shown in greater detail. The depicted images were cropped such that they have the same shape.
Figure 3: Lens contamination examples in $\mathcal{D}^{\text{folsom}}$: an image with severe contamination (left), an image taken immediately before lens cleaning (middle), and one taken immediately after cleaning (right).
@@ 3.4 Sky Image Data Processing
The image processing was essentially the same for all 3 datasets. First we compute the mean image of each dataset (using only the long exposure images for $\mathcal{D}^{\text{sirta}}$ and $\mathcal{D}^{\text{nrel}}$) and then define a circular region of interest (ROI), with radius $R$ and center $C$, which is used as a binary mask for that dataset (see Table 1). These masks are applied to all the images in their respective datasets to erase every pixel outside their ROI. This is done with the intention of standardizing the image data across all datasets, such as removing the timestamp annotation in the upper left corner of the $\mathcal{D}^{\text{nrel}}$ images and removing any lens flare and/or noise effect in the black area of the images. Following this, we center square crop the images on $C$ with length $2R$. Finally, the images are downscaled to a $64\times 64$ resolution using the Python’s PIL library. To avoid any potential JPEG compression losses [51], the resized images are saved to disk in .npy format.
Figure 4: The GHI data distributions for the three datasets used in this study, plotted after the data processing pipeline. This distribution is based on the total number of image/irradiance pairs, not the total number of irradiance measurements.
Table 1: Circular region of interest (ROI) radius and center values used in each dataset.
Dataset $R$ (pixels) $C$ (x, y)
Folsom 755 (768, 768)
SIRTA 340 (502, 394)
NREL 755 (768, 768)
@@ 3.5 Irradiance Data Processing
The irradiance data processing was also largely the same for all 3 datasets. First, we resample the data to a 1-second frequency through linear interpolation (more details on why this is done in the following section). To avoid large interpolation errors, this is only done for the samples that are 1-minute apart, i.e. periods with 2 or more minutes of missing data are discarded. Following the linear interpolation, we remove all samples where $\theta_{z}>80^{\circ}$, as these samples provide little information for our application.
The next processing step is only applied to the training set of each dataset (see subsection 3.8 for the train/test splits) and consists of shifting the irradiance measurements in time by applying a time offset ($\Delta\text{t}$), which we do according to:
@@
where $T_{\text{new}}$ and $T_{\text{orig.}}$ are the new and original irradiance measurement timestamps, respectively. If a $\Delta\text{t}<0$ is applied, then we are advancing the irradiance series in time, relative to the sky image series, whereas a $\Delta\text{t}>0$ would have the opposite effect, as shown in Figure 5. Depending on how these irradiance values were averaged, we hypothesize that this can be beneficial. For instance, if the value of $I$ at 14:00:00 represents the average GHI values from 13:59:01 to 14:00:00 (backward average), then a $\Delta\text{t}<0$ could diminish the delay introduced by such an average. Thus, we shift the irradiance measurements by -20s, 10s and 30s in $\mathcal{D}^{\text{folsom}}$, $\mathcal{D}^{\text{sirta}}$ and $\mathcal{D}^{\text{nrel}}$, respectively, but also investigate other time offsets in subsubsection 5.2.2. The reason why a $\Delta\text{t}$ is not applied to the testing set is because doing so would lead to unfair comparisons among the nowcasting results. For example, if $\Delta\text{t}>0$ (e.g. 30s) is applied to the testing set, then, given the same ASI context, the evaluation of a 10 minute lead time prediction in the shifted time series would be equivalent to a 9:30 minutes lead time in the unshifted time series, which is fundamentally easier to predict. Similarly using $\Delta\text{t}<0$ (e.g. -30s) would be equivalent to a 10:30 minute lead time, which is fundamentally harder.